OpenStudy (anonymous):

Calc 2 help please! Find the derivative of d/dx (integral) (top limit) 3 sq rt x (bottom limit) x/6 cos t dt.... I know how to solve this with standard constants in the integral but the variable/ sq rt has thrown me off.

terenzreignz (terenzreignz):

Oh really? ^_^ Then I suggest you solve it first with what you call 'standard constraints' :P Solve the coloured bit, as I ever so slightly tweaked it :D \[\Large \frac{d}{dx}\color{blue}{\int\limits_a^b\cos(t)dt}\] There... those are what you'd call standard constraints, right? XD

OpenStudy (anonymous):

well with that its just gonna be the deriv of cos t dt so cos t..

terenzreignz (terenzreignz):

Don't mind the derivative d/dx first. Just solve the blue integral.

terenzreignz (terenzreignz):

Or... you still weren't taught how to take integrals?

OpenStudy (anonymous):

the example before this we solved had the integral as t^9 upper 0 lower, and 3 sq rt u du to the right, so we simply reworked that to the derivative. so i'm confused on the integral part yeah

terenzreignz (terenzreignz):

Okay, allow me to present you a formula then ^_^ \[\Large \frac{d}{dx} \int \limits_{a(x)}^{b(x)}f(t) dt = f\left(b(x)\right)b'(x) -f\left(a(x)\right)a'(x)\]

OpenStudy (anonymous):

so input each limit on the integral times their derivative and take the difference?

terenzreignz (terenzreignz):

Yup. This is a result of the chain rule. But I suppose you'll get to that later. Before you do this, what was the answer to that other integral, the one with a t^9 upper limit?

OpenStudy (anonymous):

9t^11

terenzreignz (terenzreignz):

Are you sure it's 3sqrt? Don't you mean cube root?

OpenStudy (anonymous):

we did 3 sq rt u du -> u^ 1/3 du -> 3/4 u ^ 4/3 -> 3/4 (t^9) ^ 4/3 -> 3/4 t ^ 12 take the deriv of that and 9t^11

terenzreignz (terenzreignz):

Like this: \[\Large \sqrt[3]{\color{white}{\qquad}}\]

OpenStudy (anonymous):

yeah thats what i meant sorry

OpenStudy (anonymous):

so given we just have cos t dt i didnt know where to begin

terenzreignz (terenzreignz):

Okay, well let me just demonstrate to you that the formula I gave you works (and is actually more general) \[\Large \frac{d}{dt} \int \limits_{a(t)}^{b(t)}f(u) du = f\left(b(t)\right)b'(t) -f\left(a(t)\right)a'(t)\] ^this is the formula, tweaked in terms of u and t, right? :)

terenzreignz (terenzreignz):

Okay, in this case, our a(t) is 0 b(t) is t^9 and \[\large f(u) = \sqrt[3]u \] correct?

OpenStudy (anonymous):

yes.

terenzreignz (terenzreignz):

Johnny...! <snaps fingers> Focus! LOL I'm TJ by the way ^_^

terenzreignz (terenzreignz):

Oh, you're here... whoops :3

terenzreignz (terenzreignz):

Now, \[\Large f\left(b(t)\right)= f(t^9) = \sqrt[3]{t^9} = t^3\] Still with me?

terenzreignz (terenzreignz):

why? I just substituted...

OpenStudy (anonymous):

meant yes

terenzreignz (terenzreignz):

Right. Now, b'(t) is simply the derivative of t^9 and this is just 9t^8 Right? So we can replace this part: \[\Large \frac{d}{dt} \int \limits_{0}^{t^9}\sqrt[3]u du =\color{red}{ f\left(b(t)\right)b'(t)} -f\left(a(t)\right)a'(t)\] With this: \[\Large \frac{d}{dt} \int \limits_{0}^{t^9}\sqrt[3]u du =\color{blue}{ (t^3)(9t^8)} -f\left(a(t)\right)a'(t)\] Catch me so far?

OpenStudy (anonymous):

roger

terenzreignz (terenzreignz):

now, a(t) is simply 0 so actually, both f(a(t)) and a'(t) are going to be zero. \[\Large \frac{d}{dt} \int \limits_{0}^{t^9}\sqrt[3]u du =\color{}{ (t^3)(9t^8)} -\cancel{f\left(a(t)\right)a'(t)}^0\] So we have: \[\Large \frac{d}{dt} \int \limits_{0}^{t^9}\sqrt[3]u du =\color{}{ (t^3)(9t^8)}=9t^{11}\] And voila! It worked ^_^

OpenStudy (anonymous):

so for this question, cube root x is the upper, but we cant simplify that down, or would you have f(x^1/3)(x/3) for the first part??

OpenStudy (anonymous):

f(b(x))b'(x)= (x^1/3)(x/3)?

terenzreignz (terenzreignz):

Unfortunately :D And slow down, lol what's the derivative of the cube root of x? I don't think it's it's as simple as x/3 :P

OpenStudy (anonymous):

crap, 1/3x^2/3?

terenzreignz (terenzreignz):

\[\Large \frac{d}{dx}x^\frac13 = \frac13x^{-\frac23}\] :P

OpenStudy (anonymous):

really? wheres the negative from?

terenzreignz (terenzreignz):

Seriously? lol \[\Large \frac{d}{dx}x^a = ax^{a-1}\] And take \[\large a = \frac13\] XD

OpenStudy (anonymous):

ugh ok

terenzreignz (terenzreignz):

Convinced? Power rule, mate ^_^

OpenStudy (anonymous):

so (x^1/3)(1/3x^-2/3) -

OpenStudy (anonymous):

(x/6)(1/6)?

terenzreignz (terenzreignz):

Nope :P

OpenStudy (anonymous):

sigh

terenzreignz (terenzreignz):

Remember, it's f(b(t)) not simply b(t) :D

terenzreignz (terenzreignz):

In your case, what was f? (f was the function being integrated, look on it now :) ) \[\Large \frac{d}{dx} \int \limits_{a(x)}^{b(x)}f(t) dt = f\left(b(x)\right)b'(x) -f\left(a(x)\right)a'(x)\]

OpenStudy (anonymous):

so everything input should be multiplied by cos t?

terenzreignz (terenzreignz):

not multiplied, silly. evaluated under cosine XD

OpenStudy (anonymous):

but how the heck can we eval 1/3x^ -2/3 under cos?

terenzreignz (terenzreignz):

please study this formula more CAREFULLY \[\Large \frac{d}{dx} \int \limits_{a(x)}^{b(x)}f(t) dt = f\left(b(x)\right)b'(x) -f\left(a(x)\right)a'(x)\] What are evaluated under the f function? Are they the derivatives of a and b? NO! Just a and b themselves. And don't worry about evaluating, just express as is.

terenzreignz (terenzreignz):

If you don't like it this way, then maybe this is clearer: \[\Large \frac{d}{dx} \int \limits_{a(x)}^{b(x)}f(t) dt =[f\circ b](x)b'(x) -[f\circ a](x)a'(x)\]

OpenStudy (anonymous):

lets stick to what we had, i think i just need to see the eventual simplification of it

terenzreignz (terenzreignz):

Right... so what did you get, given my corrections...?

terenzreignz (terenzreignz):

And leave the stuff inside the cosine as is, you can't possibly evaluate them to exactness...

OpenStudy (anonymous):

so rather than x^1/3 * 1/3x^ -2/3 i should have that integrated with the cos t we started with?

OpenStudy (anonymous):

(cos^1/3)(1/3cos^-2/3) such as that?

terenzreignz (terenzreignz):

*cos x^1/3 Otherwise, yes. :)

OpenStudy (anonymous):

ok so only the first part we are using the function got it sorry

terenzreignz (terenzreignz):

So do you have a new answer? :)

OpenStudy (anonymous):

(cos x^1/3)(1/3 cos^-2/3)-

OpenStudy (anonymous):

(cos x/6)(1/6)

terenzreignz (terenzreignz):

Bingo ^_^ You can arrive at that answer another way, however, if all you had trouble with was integrating cos, then just recall that \[\large \int \cos(t) dt = \sin(t)+C\] And evaluate at the limits... and then differentiate :D

OpenStudy (anonymous):

roger

OpenStudy (anonymous):

now for the simplification of that....

terenzreignz (terenzreignz):

So basically the way to approach it was: \[\Large \frac{d}{dx} \int\limits_\frac x6 ^ \sqrt[3]x \cos(t) dt = \frac{d}{dx}\left[\sin(\sqrt[3]x ) -\sin\left(\frac x6\right)\right]\]

terenzreignz (terenzreignz):

As for simplification, it's only a few adjustments to the fractions, nothing major, and there can be no more simplification unless you know what x is.

OpenStudy (anonymous):

they have the answer as 1/3x^-2/3 cos cube root of x,...so the front half of the formula we just did

terenzreignz (terenzreignz):

Just the front half, eh? That's really interesting... what could have caused the other 'half' to zero out, I wonder... Are you sure you told us the right limits?

OpenStudy (anonymous):

cube root of x upper, lower = pi/6 , wrote x /6 down in notes for some reason

OpenStudy (anonymous):

which makes it cos pi/6 * 0 which cancels out yes, thank you though

terenzreignz (terenzreignz):

Okay, then everything's settled :P